OpenStudy (anonymous):

Given the function f(x) = 2(x + 6), find x if f(x) = 22.

OpenStudy (anonymous):

ok , plug in the value of x in the equation \[2(x+6) = 2(22+6)\] \[2(28) = 56\]

OpenStudy (anonymous):

got it gabbypriscilla

OpenStudy (anonymous):

i think it wud be 22= 2(x+6)

OpenStudy (anonymous):

22= 2x+12

OpenStudy (anonymous):

10=2x

OpenStudy (anonymous):

x=5

OpenStudy (anonymous):

becoz we r finding x

OpenStudy (anonymous):

i agree @Alisha

OpenStudy (anonymous):

oh sorry , yes yes alisha u r very correct . 22 = 2(x+6) 22 = 2x+12 22-12 =2x 10=2x 10/2 = x 5 - x actually i thought that it is f(22) , but it was f(x)= 22